Text Size

February 12, 2010
ISS On-Orbit Status 02/12/10

All ISS systems continue to function nominally, except those noted previously or below. Sleep shift in effect: The ISS crew’s workday began last evening at 4:14pm and ended this morning at 7:44 EST (see time table at bottom).

Mission 20A’s EVA-1 was completed successfully by EV1 Bob Behnken & EV2 Nicholas Patrick in 6h 32m, accomplishing all objectives plus a few get-aheads. Beginning last night at 9:17pm EST, the spacewalk ended this morning at 3:49am. [EV1 & EV2 began their “campout” yesterday evening in the U.S. Airlock (A/L) with hatch closure and depressurization of the Crewlock (CL) from 14.7 to 10.2 psi, followed by mask prebreathe. Following the usual hygiene break/with mask prebreathe for Behnken & Patrick last night, the A/L hatch was closed again by TJ Creamer for EVA preps in 10.2 psi, followed by EMU purge (~7:35pm) and prebreathe in the EMUs (7:49pm-8:39am). Afterwards, with CL depressurization and EV1/EV2 switching to suit power, EVA-1 began at 9:17pm. The excursion lasted 6h 32m.]

During EVA-1, Behnken & Patrick –
  • Prepared Node-3 “Tranquility” for unberthing,
  • Completed NH3 (Ammonia) jumper and MLI (Multi-Layer Insulation) bag activities,
  • Removed the OTP (ORU {On-orbit Replaceable Units} Temporary Platform) from the SPDM (Special Purpose Dexterous Manipulator) “Dextre”,
  • Installed the Node-3-to-ISS LTA (Launch-to-Activation) cable, required for Node-3 keep-alive heaters,
  • Installed Node-3 avionics cables, and
  • Cleaned up & ingressed.

Due to early completion of the planned EVA tasks, the spacewalkers also completed the following EVA-1 (get-ahead) tasks:
  • Relocated the NH3 Jumper bag to the deploy location,
  • Opened the Node-3 Nadir CBCS (Centerline Berthing Camera System) flap, and
  • Removed the launch tape for the ground installed gap spanners on Tranquility.

Node-3 was transferred by Kay Hire & Terry Virts on the SSRMS (Space Station Remote Manipulator System), assisted by Williams monitoring CBM (Common Berthing Mechanism), and attached to the Node-1 Port CBM. [Robotics system performance was nominal. SSRMS remains grappled to Node-3 overnight.]

Node-3 Vestibule gross leak check was completed by CDR Williams, and keep-alive heaters were activated. Fine leak checks are being conducted “overnight”.

Post-ingress activities by Behnken, Patrick, Zamka & Creamer included the usual post-EVA tasks like photographing EMU gloves for inspection, recharging EMUs with water, downloading & downlinking D2XS EVA & glove photographs, recharging REBA batteries, etc.

At wake-up last evening, before the spacewalk went underway, FE-4 Kotov did the regular daily early-morning check of the aerosol filters at the Russian Elektron O2 generator which Suraev had installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [FE-4 again inspects the filters tonight at bedtime, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

Also at wake-up, the FE-1 terminated his 9th experiment session, started last night, for the long-term Russian sleep study MBI-12/SONOKARD, by taking the recording device from his SONOKARD sports shirt pocket and later copying the measurements to the RSE-MED laptop for subsequent downlink to the ground. [SONOKARD objectives are stated to (1) study the feasibility of obtaining the maximum of data through computer processing of records obtained overnight, (2) systematically record the crewmember’s physiological functions during sleep, (3) study the feasibility of obtaining real-time crew health data. Investigators believe that contactless acquisition of cardiorespiratory data over the night period could serve as a basis for developing efficient criteria for evaluating and predicting adaptive capability of human body in long-duration space flight.]

CDR Williams & FE-6 Creamer continued their new week-long session of the experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), TJ’s second, donning their Actiwatches, from which to log data to the HRF-1 (Human Research Facility 1) laptop. [To monitor the crewmembers’ sleep/wake patterns and light exposure, the crewmembers sometimes wear a special Actiwatch device which measures the light levels encountered by them as well as their patterns of sleep and activity throughout the Expedition and use the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]

Oleg Kotov transferred the Russian VETEROK (“Breeze”) science hardware to the SM, set it up with its electrical connections, and activated it for the first checkout run, taking photography, logging parameters and downloading data to laptop. The session was supported by ground specialist tagup. [The objective was to verify operational integrity of the hardware and the efficiency of new technologies for optimizing atmospheric parameters under conditions of orbital flight. The experiment studies the implementation of alternative methods for cleaning & revitalizing the atmosphere by pumping the air with an electrostatic fan through an electric filter and saturating the airflow with light air ions of positive and negative polarity, which may solve the problem of removing organic trace contaminants from the air, both in the entire station volume and in the space behind the panels.]

In the FGB, Maxim Suraev used microbial growth wipes and Fungistat disinfectant to clean areas behind panel 206 which have shown some microbial contamination. [Areas of interest were accessible frame sections, attachments, mounting bracket, pressurized shell surface areas, panel 206 internal surface, etc.]

Continuing the installation of a new BRI Smart Switch Router computer and ASP Network Connection Adapter started yesterday, Kotov & Suraev worked in the SM, laying new cables from the BRI to the old BRI location behind the medical locker and from the new BRI to the RSS1 and RSS2 laptops, plus from the ASP to future payload Ethernet connections behind panel 225.

Maxim downloaded TEKh-15/DAKON-M IZGIB (“Bend”) structural dynamics measurements of the 20A docking and re-set the hardware for more data taking, later checking its status. [IZGIB has the objective to help update mathematical models of the ISS gravitation environment, using accelerometers of the Russian SBI Onboard Measurement System, the GIVUS high-accuracy angular rate vector gyrometer of the SUDN Motion Control & Navigation System and other accelerometers for unattended measurement of micro-accelerations at science hardware accommodation locations - (1) in operation of onboard equipment having rotating parts (gyrodynes, fans), (2) when establishing and keeping various ISS attitude modes, and (3) when performing crew egresses into space and physical exercises.]

In the Kibo JPM (JEM Pressurized Module), FE-5 Noguchi serviced the Dewey’s Forest science payload, which requires periodic watering for the cultivation of its PUs (Plant Units). In addition, Soichi rearranged the PUs so that they will have equal exposure to light. [Dewey’s Forest, one of the Japanese educational payloads, is intended to show how gravity controls the laws of nature and influences our ways of thinking. The project is “a catalyst to rediscover our relationship with plants on the ground and the age-old history of our gardens.”]

Starting today you'll be performing activities that require CSA-O2 checks behind closeout panels. You may notice some procedures have you check the labels on the CSA-O2s as a reference to acceptable O2 ranges and some procedures have you call down to MCC-H to perform the verification. We’d like to apologize up front for the mixed configuration. The work associated with labeling the CSA-O2s was only recently completed, thus only a portion of our procedures have incorporated the use of those labels. We are working to migrate all procedures to reference the labels and not require communication with MCC-H, hence for your activities, you may use the CSA-O2 labels for all verification steps.

Later, FE-4 activated & checked four newly-delivered CSA-CP (Compound Specific Analyzer-Combustion Products) units, changed out the batteries in both CSA-CP sampling pumps and the four new units (#1052, #1042, #1056, #1049), verified that their sensors are not contaminated and zero-calibrated them for use on ISS.

Noguchi conducted routine maintenance on the four CSA-CP (Compound Specific Analyzer-Combustion Products) units (#1052, #1042, #1056, #1049), changing the battery of the prime unit and performing zero calibration on all units.

Soichi also created labels and applied them to two 20A-delivered CSA-O2 units (#1041, #1045) slated for extensive upcoming atmospheric O2 analyses in the cabin and behind closeouts. [The work associated with labeling the CSA-O2s was only recently completed on the ground, thus only a portion of onboard procedures have incorporated the use of those labels. Work is underway to migrate all procedures to reference the labels, so that communication with MCC-H will not be required.]

UPA Update: Urine Processor Assembly processing was successful last night after the installation of the new unit, filling the WPA (Water Processor Assembly) waste tank, along with condensate input, to 35%. Ground engineers are now assessing low pressure indications in the WPA following the filter changeout, which could possibly slow down UPA processing desired for RFTA return. It is unlikely that the pressure drop is in the new Pump/Sep ORU, which is now protected by the EFA filter (300 microns), but the filter may already be clogged with particulates not removed from the waste tank drain. Recommended forward plan is to continue with WPA operations with the current EFA filter and to increase waste tank quantity to >65% for next process cycle to generate enough flow for MLS (Mostly Liquid Separator) cycles to occur.

The crewmembers worked out with their regular 2-hr physical exercise on the CEVIS cycle ergometer (CDR, FE-5, FE-6), TVIS treadmill (FE-1, FE-4), ARED advanced resistive exerciser (CDR, FE-1, FE-5, FE-6), and VELO bike ergometer with bungee cord load trainer (FE-4).

  • Sleep shifting started with the Progress docking on 2/4. On 2/6, crew wake shifted earlier, to 5:40pm EST. 20A Undock will drive Crew Wake one and a half hours earlier to 4:09pm by FD12. This shift is accomplished by moving Crew Sleep 60 min earlier on FDs 11-12. Wake/Sleep table:

4:14pm (2/11)
4:14pm (2/12)
4:14pm (2/13)
4:14pm (2/14)
4:14pm (2/15)
4:14pm (2/16)
4:14pm (2/17)
3:14pm (2/18)

No CEO (Crew Earth Observation) photo targets uplinked for today.

Significant Events Ahead (all dates Eastern Time and subject to change):
02/10/10 -- STS-130/Endeavour/20A docking (12:57am)
· 02/13/10 -- EVA-2 (9:14pm; Node-3 outfitting), Node-3 activation
· 02/14/10 -- Cupola relocation & outfitting
· 02/16/10 -- EVA-3 (9:14pm; Node-3, Cupola, and PMA3 tasks)
02/17/10 -- Final cargo transfers, hatch closure, reboost
02/18/10 -- STS-130/Endeavour/20A undock (7:35pm)
02/20/10 -- STS-130/Endeavour/20A KSC deorbit burn (8:59pm)
02/20/10 -- STS-130/Endeavour/20A KSC landing (10:01pm)
03/18/10 -- Soyuz TMA-16/20S undock/landing
--------------Three-crew operations-------------
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC (launch ~1:30pm EST)
04/02/10 -- Soyuz TMA-18/22S launch – Skvortsov (CDR-24)/Caldwell/Kornienko
04/04/10 -- Soyuz TMA-18/22S docking
--------------Six-crew operations-----------------
04/27/10 -- Progress M-03M/35P undock
04/28/10 -- Progress M-05M/37P launch
04/30/10 -- Progress M-05M/37P docking
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1 (~2:00pm EST)
05/10/10 -- Progress M-04M/36P undock
05/31/10 -- Soyuz TMA-17/21S undock/landing
--------------Three-crew operations-------------
06/14/10 -- Soyuz TMA-19/23S launch – Wheelock (CDR-25)/Walker/Yurchikhin
06/16/10 -- Soyuz TMA-19/23S docking
--------------Six-crew operations-----------------
07/xx/10 -- US EVA-15
07/xx/10 -- Russian EVA-25
06/28/10 -- Progress M-06M/38P launch
07/02/10 -- Progress M-06M/38P docking
07/26/10 -- Progress M-05M/37P undock
07/27/10 -- Progress M-07M/39P launch
07/29/10 -- Progress M-07M/39P docking
07/29/10 -- STS-134/Endeavour (ULF6 – ELC3, AMS-02) (~7:30am EST)
08/30/10 -- Progress M-06M/38P undock
08/31/10 -- Progress M-08M/40P launch
09/02/10 -- Progress M-08M/40P docking
09/15/10 -- Soyuz TMA-18/22S undock/landing
09/16/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) (~12:01pm EST)
09/18/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) docking
09/22/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) undock
09/30/10 -- Soyuz TMA-20/24S launch – Kelly (CDR-26)/Kaleri/Skripochka
10/xx/10 -- Russian EVA-26
10/26/10 -- Progress M-07M/39P undock
10/27/10 -- Progress M-09M/41P launch
10/29/10 -- Progress M-09M/41P docking
11/15/10 -- Soyuz TMA-19/23S undock/landing
11/30/10 -- Soyuz TMA-21/25S launch – Kondratyev (CDR-27)/Coleman/Nespoli
12/15/10 -- Progress M-08M/40P undock
02/08/11 -- Progress M-09M/41P undock
02/09/11 -- Progress M-10M/42P launch
02/11/11 -- Progress M-10M/42P docking
03/30/11 -- Soyuz TMA-22/26S launch
xx/xx/11 -- Progress M-11M/43P launch
05/30/11 -- Soyuz TMA-23/27S launch
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.